# Multiscale Biomolecular Simulations in the Exascale Era

### v1

@@ 5 Outlook
The versatility of the MiMiC framework makes it ideal for pushing multiscale biomolecular simulations towards the exascale. It is capable of exploiting the results of the large-scale initiatives by EuroHPC JU and ECP for individual domains such as QM and MM. Indeed, work is already underway to couple a diverse set of QM programs to MiMiC, namely, Quantum ESPRESSO, CP2K, and DFT-FE, all of which are being developed to exploit state-of-the-art HPC technology [59]. Additionally, it is likely that MD simulations will benefit greatly from future ML-based force fields [68, 69], which MiMiC is fully able to integrate into advanced simulation workflows [70]. For instance, hybrid ML/MM models facilitate long simulations of biological systems using ML-based force fields, achieving near QM/MM precision at a substantially reduced computational cost [71], while ML-enhanced free-energy methods significantly accelerate QM/MM calculations of ligand binding affinities for drug discovery [72]. These and other novel methods are expected to aid in accurately predicting key biophysical properties like drug-protein binding free energies and complete free-energy profiles. We anticipate that the integration of MiMiC-based multiscale simulations with ML and enhanced sampling techniques will further catalyze a quantum leap in the fundamental atomistic understanding of biological processes.
@@ 5.1 Acknowledgment
This preprint was created using the LaPreprint template (https://github.com/roaldarbol/lapreprint) by Mikkel Roald-Arbøl .

### v2

@@ 5 Outlook
The versatility of the MiMiC framework makes it ideal for pushing multiscale biomolecular simulations towards the exascale. It is capable of exploiting the results of the large-scale initiatives by EuroHPC JU and ECP for individual domains such as QM and MM. Indeed, work is already underway to couple a diverse set of QM programs to MiMiC, namely, Quantum ESPRESSO, CP2K, and DFT-FE, all of which are being developed to exploit state-of-the-art HPC technology Gavini2023. Additionally, it is likely that MD simulations will benefit greatly from future ML-based force fields Unke2021; Kocer2022, which MiMiC is fully able to integrate into advanced simulation workflows Mouvet2022. For instance, hybrid ML/MM models facilitate long simulations of biological systems using ML-based force fields, achieving near QM/MM precision at a substantially reduced computational cost Fabritiis, while ML-enhanced free-energy methods significantly accelerate QM/MM calculations of ligand binding affinities for drug discovery Rizzi2023. These and other novel methods are expected to aid in accurately predicting key biophysical properties like drug-protein binding free energies and complete free-energy profiles. We anticipate that the integration of MiMiC-based multiscale simulations with ML and enhanced sampling techniques will further catalyze a quantum leap in the fundamental atomistic understanding of biological processes.
@@ 5.1 Acknowledgment
[…]